if x = a cos t , y = b sin t . Then find the value of d^3y/dx^3 = ?

 if  x = a cos t , y = b sin t . Then find the value of d^3y/dx^3 = ?
 

Grade:11

1 Answers

Khimraj
5 years ago
x = acost
dx/dt = -asint
d2x/dt2 = - acost
d3x/dt3 = asint
y = bsint
dy/dt = bcost
d2y/dt2 = – bsint
d3y/dt3 = – bcost  
d3y/dx3 = (d3y/dt3)/(d3x/dt3) = asint/(-bcost) = -(a/b)tant
Hope it clears. If you like answer then please approve it.